Neil Armstrong was the first man to walk on the moon. On July 20, 1969, his famous statement, "That's one small step for man, one giant leap for mankind." was received at 10:56 p.m. ( EDT) on Earth. (a) How long did it take his voice to reach Earth via radio waves? (b) At what time (in EDT) did he utter this statement at the surface of the moon. The distance between the Earth and the moon is 3.85 x 10^12 m.



1
Expert's answer
2022-04-19T12:05:51-0400

First, note, that the given distance from Earth to Moon is not correct. The right value is 3.85 x 10^8 m

a) The time required for the light to reach Earth from Moon is:


"t = \dfrac{d}{c}"

where "d = 3.85 \times10^{8} m" and "c = 3.00\times 10^8m/s" is the speed of light. Thus, obtain:


"t = \dfrac{3.85 \times10^{8}m}{ 3.00\times 10^8m/s} \approx 1.28s"

b) Thus, the time he utter this statement at the surface of the moon is approximately 10:55 p.m.